# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v2

@@ 2.4.2 Decoding Configuration
1 | PolyCoder | 0.4B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 273 | 1.8K |
2 | InCoder | 1.3B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 3060 | 3.2K |
3 | Phi 1 | 1.3B | Decoder | MIT | Jun 2023 | Microsoft | 7449 | N/A |
4 | DeepSeek-Coder | 1.3B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 123488 | 21K |
5 | OpenCodeInterpreter | 1.3B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 555 | 1.6K |
6 | Yi-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 8913 | 7.8K |
7 | Qwen2.5-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 11799 | 4.6K |
@@
8 | CodeGemma | 2.0B | Decoder | Gemma | Apr 2024 | 37008 | N/A | |
9 | PolyCoder | 2.7B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 330 | 1.8K |
10 | Stable Code | 3.0B | Decoder | Apache 2.0 | Jan 2024 | Stability AI | 175 | 125 |
11 | StarCoder2 | 3.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 1468288 | 1.9K |
12 | Qwen2.5-Coder | 3.0B | Decoder | Qwen Research | Sep 2024 | Alibaba Cloud | 13540 | 4.6K |
@@
13 | InCoder | 6.7B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 640 | 3.0K |
14 | DeepSeek-Coder | 6.7B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 28065 | 21K |
15 | OpenCodeInterpreter | 6.7B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 136 | 1.5K |
16 | CodeLlama | 7.0B | Decoder | Llama2 | May 2024 | Meta | 180516 | 16.2K |
17 | StarCoder2 | 7.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 125386 | 2.0K |
18 | CodeGemma | 7.0B | Decoder | Gemma | Apr 2024 | 42304 | N/A | |
19 | Qwen2.5-Coder | 7.0B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 64152 | 4.6K |
20 | Yi-Coder | 9.0B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 49711 | 7.8K |
@@ 3.2 Performance Trade-Offs (RQ2)
1 | PolyCoder_0.4B | 0.01 | 0.02 | 0.02 | 0.02 | 5.03 | 2.49 |
2 | Incoder_1.3B | 0.06 | 0.08 | 0.11 | 0.13 | 13.88 | 3.06 |
3 | Phi1_1.3B | 0.41 | 0.45 | 0.49 | 0.50 | 8.45 | 1.28 |
4 | DeepSeek-Coder_1.3B | 0.43 | 0.48 | 0.54 | 0.57 | 12.12 | 1.19 |
5 | OpenCodeInterpreter_1.3B | 0.53 | 0.58 | 0.62 | 0.64 | 11.97 | 1.06 |
6 | Yi-Coder_1.5B | 0.49 | 0.54 | 0.58 | 0.61 | 11.99 | 0.80 |
7 | Qwen2.5-Coder_1.5B | 0.54 | 0.60 | 0.65 | 0.68 | 6.55 | 1.39 |
@@
8 | CodeGemma_2.0B | 0.28 | 0.34 | 0.39 | 0.42 | 8.70 | 1.78 |
9 | PolyCoder_2.7B | 0.03 | 0.04 | 0.05 | 0.07 | 14.15 | 5.58 |
10 | Stable Code_3.0B | 0.27 | 0.34 | 0.42 | 0.46 | 17.57 | 1.64 |
11 | StarCoder2_3.0B | 0.42 | 0.48 | 0.53 | 0.57 | 7.99 | 1.03 |
12 | Qwen2.5-Coder_3.0B | 0.59 | 0.64 | 0.69 | 0.72 | 10.77 | 1.69 |
@@
13 | Incoder_6.7B | 0.10 | 0.13 | 0.17 | 0.19 | 23.53 | 7.74 |
14 | DeepSeek-Coder_6.7B | 0.56 | 0.62 | 0.67 | 0.70 | 23.30 | 3.34 |
15 | OpenCodeInterpreter_6.7B | 0.67 | 0.70 | 0.73 | 0.74 | 23.32 | 2.67 |
16 | CodeLlama_7.0B | 0.37 | 0.44 | 0.51 | 0.55 | 23.58 | 2.11 |
17 | StarCoder2_7.0B | 0.47 | 0.52 | 0.57 | 0.60 | 17.00 | 2.92 |
18 | CodeGemma_7.0B | 0.49 | 0.59 | 0.68 | 0.74 | 23.38 | 1.67 |
19 | Qwen2.5-Coder_7.0B | 0.65 | 0.70 | 0.74 | 0.77 | 23.70 | 0.97 |
20 | Yi-Coder_9.0B | 0.59 | 0.65 | 0.71 | 0.73 | 20.33 | 3.99 |

### v3

@@ 2.4.2 Decoding Configuration
| 1 | PolyCoder | 0.4B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 273 | 1.8K |
| 2 | InCoder | 1.3B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 3060 | 3.2K |
| 3 | Phi 1 | 1.3B | Decoder | MIT | Jun 2023 | Microsoft | 7449 | N/A |
| 4 | DeepSeek-Coder | 1.3B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 123488 | 21K |
| 5 | OpenCodeInterpreter | 1.3B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 555 | 1.6K |
| 6 | Yi-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 8913 | 7.8K |
| 7 | Qwen2.5-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 11799 | 4.6K |
@@
| 8 | CodeGemma | 2.0B | Decoder | Gemma | Apr 2024 | 37008 | N/A | |
| 9 | PolyCoder | 2.7B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 330 | 1.8K |
| 10 | Stable Code | 3.0B | Decoder | Apache 2.0 | Jan 2024 | Stability AI | 175 | 125 |
| 11 | StarCoder2 | 3.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 1468288 | 1.9K |
| 12 | Qwen2.5-Coder | 3.0B | Decoder | Qwen Research | Sep 2024 | Alibaba Cloud | 13540 | 4.6K |
@@
| 13 | InCoder | 6.7B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 640 | 3.0K |
| 14 | DeepSeek-Coder | 6.7B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 28065 | 21K |
| 15 | OpenCodeInterpreter | 6.7B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 136 | 1.5K |
| 16 | CodeLlama | 7.0B | Decoder | Llama2 | May 2024 | Meta | 180516 | 16.2K |
| 17 | StarCoder2 | 7.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 125386 | 2.0K |
| 18 | CodeGemma | 7.0B | Decoder | Gemma | Apr 2024 | 42304 | N/A | |
| 19 | Qwen2.5-Coder | 7.0B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 64152 | 4.6K |
| 20 | Yi-Coder | 9.0B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 49711 | 7.8K |
@@ 3.2 Performance Trade-Offs (RQ2)
| 1 | PolyCoder_0.4B | 0.01 | 0.02 | 0.02 | 0.02 | 5.03 | 2.49 |
| 2 | Incoder_1.3B | 0.06 | 0.08 | 0.11 | 0.13 | 13.88 | 3.06 |
| 3 | Phi1_1.3B | 0.41 | 0.45 | 0.49 | 0.50 | 8.45 | 1.28 |
| 4 | DeepSeek-Coder_1.3B | 0.43 | 0.48 | 0.54 | 0.57 | 12.12 | 1.19 |
| 5 | OpenCodeInterpreter_1.3B | 0.53 | 0.58 | 0.62 | 0.64 | 11.97 | 1.06 |
| 6 | Yi-Coder_1.5B | 0.49 | 0.54 | 0.58 | 0.61 | 11.99 | 0.80 |
| 7 | Qwen2.5-Coder_1.5B | 0.54 | 0.60 | 0.65 | 0.68 | 6.55 | 1.39 |
@@
| 8 | CodeGemma_2.0B | 0.28 | 0.34 | 0.39 | 0.42 | 8.70 | 1.78 |
| 9 | PolyCoder_2.7B | 0.03 | 0.04 | 0.05 | 0.07 | 14.15 | 5.58 |
| 10 | Stable Code_3.0B | 0.27 | 0.34 | 0.42 | 0.46 | 17.57 | 1.64 |
| 11 | StarCoder2_3.0B | 0.42 | 0.48 | 0.53 | 0.57 | 7.99 | 1.03 |
| 12 | Qwen2.5-Coder_3.0B | 0.59 | 0.64 | 0.69 | 0.72 | 10.77 | 1.69 |
@@
| 13 | Incoder_6.7B | 0.10 | 0.13 | 0.17 | 0.19 | 23.53 | 7.74 |
| 14 | DeepSeek-Coder_6.7B | 0.56 | 0.62 | 0.67 | 0.70 | 23.30 | 3.34 |
| 15 | OpenCodeInterpreter_6.7B | 0.67 | 0.70 | 0.73 | 0.74 | 23.32 | 2.67 |
| 16 | CodeLlama_7.0B | 0.37 | 0.44 | 0.51 | 0.55 | 23.58 | 2.11 |
| 17 | StarCoder2_7.0B | 0.47 | 0.52 | 0.57 | 0.60 | 17.00 | 2.92 |
| 18 | CodeGemma_7.0B | 0.49 | 0.59 | 0.68 | 0.74 | 23.38 | 1.67 |
| 19 | Qwen2.5-Coder_7.0B | 0.65 | 0.70 | 0.74 | 0.77 | 23.70 | 0.97 |
| 20 | Yi-Coder_9.0B | 0.59 | 0.65 | 0.71 | 0.73 | 20.33 | 3.99 |